### v1

# DomainDemo: a dataset of domain-sharing activities among different demographic groups on Twitter

Abstract
Social media play a pivotal role in disseminating web content, particularly during elections, yet our understanding of the association between demographic factors and political discourse online remains limited. Here, we introduce a unique dataset, DomainDemo, linking domains shared on Twitter (X) with the demographic characteristics of associated users, including age, gender, race, political affiliation, and geolocation, from 2011 to 2022. This new resource was derived from a panel of over 1.5 million Twitter users matched against their U.S. voter registration records, facilitating a better understanding of a decade of information flows on one of the most prominent social media platforms and trends in political and public discourse among registered U.S. voters from different sociodemographic groups. By aggregating user demographic information onto the domains, we derive five metrics that provide critical insights into over 129,000 websites. In particular, the localness and partisan audience metrics quantify the domains’ geographical reach and ideological orientation, respectively. These metrics show substantial agreement with existing classifications, suggesting the effectiveness and reliability of DomainDemo’s approach.
Background & Summary
Social media play a significant role in the distribution of information (Kietzmann et al., 2011; Osatuyi, 2013; Gaál et al., 2015), serving as essential sources of information for millions of users, especially in critical contexts such as elections (Himelboim et al., 2013; Holt et al., 2013) and public health crises (Gui et al., 2017; Kudchadkar and Carroll, 2020; Mitchell et al., 2020; Mitchell and Liedke, 2021). The content shared on social media originates not only from the users themselves but also from a wide array of sources. In particular, posting and re-sharing links to external websites (URLs) are key mechanisms for disseminating web content on social media (Kwak et al., 2010; Ju et al., 2014; Liedke and Wang, 2023; Rosenstiel et al., 2015). For some websites, this is a crucial way of getting traffic and for users to access their content (Cordrey, 2012; Phillips, 2012; Newman et al., 2016; Myllylahti, 2018). For instance, Chen et al. (2022) demonstrate that social media are responsible for a substantial portion of referred visits to thegatewaypundit.com, a popular far-right news website, by analyzing its web traffic data.
The significance of social media in information distribution has attracted considerable research attention in recent years. However, our understanding of the role demographic factors play in this process remains limited (Golder and Macy, 2014; Shugars et al., 2021), even though their importance in online discourse has been discussed in prior work (Allen et al., 2013; Ting and Ahmad, 2017; Falenska et al., 2024; Barberá and Rivero, 2015). Typically, the data provided by platforms to researchers lack user-level demographic information (Veale and Binns, 2017; Holstein et al., 2019; Andrus et al., 2021). When researchers rely on user donations or surveys to collect demographic data, the sample size is often insufficient to provide meaningful aggregate insights about content-sharing patterns, particularly at the domain level. These challenges in data collection have created a significant gap in our ability to comprehensively analyze the interplay between demographic factors and online political discourse.
To bridge this gap, we introduce a novel dataset, DomainDemo, which quantifies domain-sharing activities across diverse demographic groups on Twitter. Although Twitter was rebranded as X in 2023, we will refer to the platform as Twitter in this article since our dataset predates the rebranding. Our data encompasses demographic details such as age, gender, race, political affiliation, and geolocation (U.S. state). These domain-sharing events are derived from a comprehensive dataset of over 1.5 million U.S.-based Twitter users matched with their voter registration records. Spanning 132 months (11 years) from May 2011 to April 2022, the dataset is organized in monthly intervals, enabling the analysis of temporal trends. Our released datasets allow researchers to investigate the association between demographic characteristics and the sharing patterns of diverse information sources, ranging from mainstream news websites to potential sources of misinformation.
We release two variants of the domain-sharing statistics dataset: DomainDemo-multivariate and DomainDemo-univariate. The DomainDemo-multivariate variant splits the statistics into buckets defined by domain, age, gender, race, political affiliation, state, and month altogether. In each bucket, we provide the number of shares, the number of unique users sharing the domain, and the Gini index (Gini, 1912) of the sharing count among users. The DomainDemo-univariate variant includes five universes: state, race, gender, age, and political affiliation. Within each universe, we provide the number of shares, the number of unique users sharing the domains, and the Gini index of the sharing count among users for each category (e.g., age groups in the age universe).
The DomainDemo-multivariate variant is similar in format to the Facebook URL dataset shared through Social Science One (King and Persily, 2020; Messing et al., 2020), which includes the number of views, shares, and reactions to various URLs across different demographic groups. However, there are key differences. DomainDemo-multivariate focuses solely on a sample of Twitter users who are registered voters in the U.S., while the Facebook URL dataset includes data from all eligible users on the platform. Additionally, DomainDemo-multivariate only includes the number of shares and unique users at the domain level and does not incorporate any noise.
In addition to the count statistics, we introduce five derived metrics that quantify how the demographics (age, gender, race, political affiliation, and geolocation) of users sharing a particular domain differ from that of the baseline. These metrics have specific interpretations that are useful for various research questions. For instance, the geolocation of the users allows us to measure the domains’ localness, i.e., the extent to which the sharing of a particular domain is geographically localized. Together with the user-sharing behavior data, the localness metric enables researchers to quantify the changing landscape of the local news industry. As a fundamental component of the U.S. democratic process, local news is uniquely positioned to report on local affairs and elections (Martin and McCrain, 2019; Moskowitz, 2021), but faces a declining trend over the years (Hayes and Lawless, 2018; Abernathy, 2018). Similarly, the user party affiliation in DomainDemo allows us to measure the audience partisanship of different domains. This metric can serve as a proxy for the political leaning of the domains, crucial for understanding online political discourse (Bakshy et al., 2015; Robertson et al., 2023).
Our derived metrics demonstrate strong alignment with established measures of localness and political leaning while significantly expanding coverage to over 129,000 domains—over ten times the number of domains in existing datasets. The metrics also uncover subtle variations in sharing patterns that previous binary or one-dimensional categorization schemes could not capture. Furthermore, our dataset enables researchers to analyze how domain-sharing behaviors have evolved across an 11-year period.
Methods
In this section, we describe how we create our dataset.
Twitter panel
Our dataset is based on a panel of over 1.5 million Twitter users matched with U.S. voter registration records. We leverage the methodology outlined in previous studies (Grinberg et al., 2019; Shugars et al., 2021; Hughes et al., 2021), utilizing the Twitter Decahose, a 10% random sample of all tweets, to identify 290 million profiles that post content between January 2014 and March 2017. These profiles are then matched with voter records provided by TargetSmart in October 2017, covering all 50 U.S. states and the District of Columbia. A Twitter account is considered a match to a voter record if the full name exactly matches and the individual is the only person with that name in the specified city or state-level geographic area in both datasets. This reliance on full names and disclosed locations helps to eliminate many fake, automated (bot), and organizational accounts.
Matching to voter file records provides access to the geolocation, year of birth, gender, race/ethnicity, and partisanship of the matched Twitter users. For geolocation, we consider only the state of residence as recorded in the voter files. Using the year of birth, we determine the age of users at the time of sharing events and categorize them into the following age groups: “<18,” “18-29,” “30-49,” “50-64,” and “>65.” The category for users younger than 18 years old is included because they might post content before reaching the legal voting age. Gender is a binary measure provided by TargetSmart, which does not capture gender identities beyond the binary framework (Shugars et al., 2024). Race/ethnicity information is inferred by TargetSmart for most states and is categorized as “African-American,” “Asian,” “Hispanic,” “Caucasian,” and “Other.”
TargetSmart provides two measures of partisanship: party registration and inferred partisanship. Party registration information in voter files is self-reported and aligns well with survey self-reporting (Ansolabehere and Hersh, 2017). However, this information is unavailable for 20 states (AL, AR, GA, HI, IL, IN, MI, MN, MO, MS, MT, ND, OH, SC, TN, TX, VA, VT, WA, and WI) in the TargetSmart data, which account for 42.7% of the Twitter users in our panel. When categorizing party registration information, we treat values for users in the 20 aforementioned states as missing. For other states, users registered as “Democrat” and “Republican” are coded accordingly. Due to variations in the classification of independent registered voters by state, we group individuals listed as “Independent,” “No party,” or “Unaffiliated” into a single “Independent” category. Members of minor parties, such as the Green Party and the Libertarian Party, are categorized as “Other.” Based on party registration and other indicators, TargetSmart infers the probability of each person voting Democrat. An evaluation from the Pew Research Center suggests this inferred partisanship is reasonably accurate (Igielnik et al., 2018). We categorize individuals as Republican (0-0.35), Independent (0.35-0.65), and Democrat (0.65-1) using TargetSmart’s recommended thresholds to generate the inferred partisanship. For our data release and analysis, we use inferred partisanship as the primary measure since it covers all users (referred to as “party” hereafter). Additionally, we provide party registration information as a secondary measure (referred to as “party registration” hereafter), as it offers useful insights for certain analyses.
Missing values in all dimensions are coded as “Unknown.”
Domain sharing statistics
We collect posts from users in the panel spanning from May 2011 to April 2022. We extract the links shared by these users, expand the shortened links when possible, and identify the corresponding domains (e.g., nytimes.com for The New York Times). This process allows us to determine which user shares which domain and when. Sharing events, as defined in our study, include posting links in original tweets and retweeting or quoting tweets containing links. To reduce noise and the risks of re-identification, we include only domains that are shared by at least 50 unique users throughout the entire period.
We integrate the demographic information of users with their domain-sharing records to construct a comprehensive table. This domain-sharing event table includes the following columns: user_id, domain, age, gender, race, party, party registration, state, and year-month. Each row corresponds to a single sharing event, with users who share the same domains multiple times contributing multiple rows. Due to the presence of user identifiers, we cannot release this detailed table. Instead, we provide aggregate statistics derived from this table, as detailed below.
To produce DomainDemo-multivariate, we group the domain-sharing event table by domain, age, gender, race, party (party registration is excluded here), state, and year-month, and calculate the following statistics: the number of shares, the number of unique users who share the domain, and the Gini index of the sharing count across users. Formally, the Gini index $G$ for a domain is calculated as:
$G$ $=\frac{2}{N^{2}\bar{x}}\sum_{i=1}^{N}\sum_{j=1}^{N}|x_{i}-x_{j}|\text{,}$ (1)
where $N$ is the number of users who share the domain, $x_{i}$ is the number of shares by the $i$th user, and $\bar{x}=\sum_{i=1}^{N}x_{i}/N$ is the average number of shares per user. $G$ ranges from 0 to 1, with 0 indicating equal sharing and values close to 1 indicating that a few users share the domain disproportionately. Note that we set the Gini index to 1 for domains shared by only one user in the bucket. We include the Gini index to help researchers understand the inequality of sharing events across users without releasing detailed information about these users due to privacy concerns.
DomainDemo-univariate is generated by aggregating the sharing events across all demographic dimensions except for the one of interest. For example, the state universe is produced by aggregating the sharing events across all age, gender, race, and party dimensions, resulting in the statistics of sharing events within different states. Similar to DomainDemo-multivariate, in each bucket of DomainDemo-univariate, we provide three statistics: the number of shares, the number of unique users who share the domain, and the Gini index of the sharing count across users.
The detailed data schema of the released versions of DomainDemo-multivariate and DomainDemo-univariate can be found in the Data Records section.
Derived metrics for domains
Based on the sharing behavior of users in different demographic groups, we derive additional metrics that quantify different aspects of the domains.
Demographic dimensions with multiple categories
Consider a demographic dimension with multiple categories indexed by $s$, such as states for geolocation. We first calculate the user distribution across all categories for all domains as the baseline distribution. We then compare the user distribution of individual domains in each category with the baseline distribution to measure the deviation. For geolocation, the deviation can be interpreted as the localness of the domain. This method can be applied to other demographic dimensions with multiple categories, such as age and race.
For a formal definition, let $C_{s}$ represent the number of unique users in category $s$ across all domains, and $F_{s}$ represent the corresponding frequency, where $F_{s}=C_{s}/\sum_{s}C_{s}$. For a domain $\delta$, we calculate the user frequency in category $s$, $F_{\delta,s}=C_{\delta,s}/\sum_{s}C_{\delta,s}$, where $C_{\delta,s}$ represents the number of unique users in category $s$ who share the domain $\delta$. For domains shared by diverse users, the observed distribution $F_{\delta,s}$ should closely align with the baseline distribution $F_{s}$ across different categories. However, deviations from the baseline distribution are expected for domains with a more concentrated audience.
Following this intuition, we quantify the deviation of a domain, denoted by $\mathcal{L}_{\delta}$, using the Kullback–Leibler (KL) divergence between $F_{\delta,s}$ and $F_{s}$:
$\mathcal{L}_{\delta}$ $=D^{(KL)}(F_{\delta,s}||F_{s})=\sum_{s}F_{\delta,s}\log_{2}\frac{F_{\delta,s}}{F_{s}}\text{,}$ (2)
where $F_{\delta,s}\log_{2}(F_{\delta,s}/F_{s})$ measures the discrepancies between the observed sharing patterns and the baseline distribution of domain $\delta$ in category $s$. $\mathcal{L}_{\delta}$ is a non-negative value that is minimized at zero when $F_{\delta,s}$ and $F_{s}$ are identical.
A limitation of $\mathcal{L}_{\delta}$ is that it can only indicate the deviation of a domain’s sharing pattern from the baseline distribution. To reveal which categories are over-represented or under-represented, one needs to further inspect the values of $F_{\delta,s}$ and $F_{s}$.
Demographic dimensions with two categories
Demographic dimensions with two categories are simpler to analyze, allowing us to directly measure the audience leaning of a domain based on the number of users from each category. For instance, in the party dimension, excluding Independent users leaves us with Democrat and Republican users. We can then assess the audience partisanship of a domain by the number of users from each party. This approach is also applicable to the gender dimension.
Formally, the audience leaning score $\mathcal{P}_{\delta}$ of a domain $\delta$ is calculated as follows:
$\mathcal{P}_{\delta}$ $=\frac{\frac{C_{\delta,R}}{C_{R}}-\frac{C_{\delta,L}}{C_{L}}}{\frac{C_{\delta,R}}{C_{R}}+\frac{C_{\delta,L}}{C_{L}}}\text{,}$ (3)
where $C_{\delta,L}$ and $C_{\delta,R}$ represent the number of unique users from the $L$ and $R$ categories who share the domain $\delta$, respectively. $C_{L}$ and $C_{R}$ represent the total number of unique users in the $L$ and $R$ categories who share any domain, respectively. Since a user can share multiple domains, we have $C_{\delta,L}\leq C_{L}\leq\sum_{i\in I}C_{i,L}$ and $C_{\delta,R}\leq C_{R}\leq\sum_{i\in I}C_{i,R}$. $\mathcal{P}_{\delta}$ is a continuous value between $-1$ and $+1$, where $-1$ means the domain $\delta$ is exclusively shared by users in the $L$ category and $+1$ means $\delta$ is exclusively shared by $R$ category users.
Data Records
Domain sharing statistics
Table 1: Schema of the DomainDemo-multivariate tables.
Column Data type Category Note
domain string key second-level domains of websites, e.g., nytimes.com
state string demographic two-letter aberrations, e.g., MA; including 50 U.S. states and the District of Columbia
race string demographic can be one of “African-American,” “Caucasian,” “Hispanic,” “Asian,”“Other,” and “Unknown”
gender string demographic can be one of “Male,” “Female,” and “Unknown”
age string demographic age bucket; can be one of “<18,” “18-29,” “30-49,” “50-64,” “65+,” and “Unknown”
party string demographic can be one of “Democrat,” “Independent,” and “Republican”
time string time year and month in the format “YYYY-MM”
shares integer statistic sharing event count of the domain in the bucket
users integer statistic unique number of users who share the domain in the bucket
gini float statistic Gini index of the sharing count across users in the bucket
We release two variants of the domain-sharing statistics dataset: DomainDemo-multivariate and DomainDemo-univariate. The table schema for DomainDemo-multivariate is provided in Table 1. For example, a row with the following values: domain=example.com, state=CA, race=Asian, gender=Female, age=30-49, party=Democrat, year_month=2018-12, shares=50, users=10, gini=0.1 indicates that the domain example.com was shared 50 times in December 2018 by 10 users who live in California, are Asian females aged between 30 and 49, and identify as Democrats. The Gini index of 0.1 indicates that the sharing count is almost evenly distributed across these users. DomainDemo-multivariate has 113,835,160 rows in total, with 129,127 unique domains. The dataset is divided into 132 parts, each corresponding to a month between May 2011 and April 2022. Note that party registration is not included in the multivariate variant.
Table 2: Schema of the DomainDemo-univariate tables. There are six universes: state, race, gender, age, party, and party registration, which share some common columns.
Data universe Column Data type Category Note
Shared by all universes domain string key second-level domains of the websites, e.g., nytimes.com
time string time year and month in the format “YYYY-MM”
shares integer statistic sharing event count of the domain in the bucket
users integer statistic unique number of users who share the domain in the bucket
gini float statistic Gini index of the sharing count across users in the bucket
state state string demographic two-letter aberrations, e.g., MA; including 50 U.S. states and the District of Columbia
race race string demographic can be one of “African-American,” “Caucasian,” “Hispanic,” “Asian,”“Other,” and “Unknown”
gender gender string demographic can be one of “Male,” “Female,” and “Unknown”
age age string demographic age bucket; can be one of “<18,” “18-29,” “30-49,” “50-64,” “65+,” and “Unknown”
party party string demographic inferred party; can be one of “Democrat,” “Independent,” and “Republican”
party registration partyreg string demographic party registration; can be one of “Democrat,” “Independent,” “Republican,” “Other,” and “Unknown”
DomainDemo-univariate details the statistics of sharing events across various categories within individual demographic variables. It includes separate sets of statistics (universes) for state, race, gender, age, party, and party registration. The table schema for these statistics is provided in Table 2. For example, a row in the state universe with the values: domain=example.com, year_month=2018-04, shares=3,000, users=250, gini=0.8, and state=NY indicates that the domain example.com was shared 3,000 times in April 2018 by 250 users in New York. The Gini index of 0.8 suggests that the sharing count is highly concentrated among a few users. Similar to DomainDemo-multivariate, each dataset within DomainDemo-univariate is divided into 132 monthly files.
Derived metrics for domains
In addition to the domain-sharing statistics, we also release the derived metrics for domains. Note that the metrics can be computed at various temporal granularities, such as monthly intervals. For the released version, we calculate these metrics over the entire 11-year period. We provide the code to enable researchers to reproduce the metrics at different temporal granularities.
Domain localness
Local news is a fundamental component of the U.S. democratic process. It is uniquely positioned to report on local affairs and elections, enabling citizens to engage in local political activities and hold their elected officials accountable (Martin and McCrain, 2019; Moskowitz, 2021). However, the landscape of local journalism is undergoing significant changes, marked by a notable decline in local agencies, often referred to as the emergence of the “news desert” (Hayes and Lawless, 2018; Abernathy, 2018). This trend threatens the vibrancy of local political participation and raises concerns about the overall health of democracy (Rubado and Jennings, 2020; Shaker, 2014).
To empirically understand the dynamics of news consumption and related phenomena, it is essential to reliably categorize news outlets as either local or national. Despite extensive research efforts, a universally accepted definition of local news organizations remains elusive (Hagar et al., 2020). Many studies on local news often fail to provide a clear definition or specific criteria for classification (Lowrey et al., 2008), complicating efforts to expand the scope of classification and hindering the replication of analyses.
Table 3: Derived metrics for domains. The table presents six metrics derived from six different data universes. From left to right, we report the metric name, size, range, data universe, calculation method, and a brief description.
Metric name Size Range Data universe Calculation Note
Localness 129,127 $[0,+\infty)$ state Eq. (2) Larger values indicate more local domains.
Race deviation 129,127 $[0,+\infty)$ race Eq. (2) Larger values indicate that the domain shares are more concentrated on some race categories.
Age deviation 129,127 $[0,+\infty)$ age Eq. (2) Larger values indicate that the domain shares are more concentrated on some age categories.
Audience partisanship 129,127 $[-1,+1]$ party Eq. (3) Only including Democrat and Republican categories. Negative values indicate more shares from Democratic users and vice versa.
Audience partisanship (registration) 129,041 $[-1,+1]$ party registration Eq. (3) Only including Democrat and Republican categories. Negative values indicate more shares from Democratic users and vice versa.
Gender leaning 129,127 $[-1,+1]$ gender Eq. (3) Only including Male and Female categories. Negative values indicate more shares from Male users and vice versa.
We leverage the state universe data from DomainDemo-univariate to derive a data-driven metric that quantifies the “localness” of news domains. This is achieved by calculating the deviation of the user distribution of each domain in different states from the baseline distribution using Eq. (2). Details of the metric are provided in Table 3.
Domain audience partisanship
A healthy democratic society requires the public to receive accurate and unbiased news and civic information, especially during election seasons (Sunstein, 2006). However, the presence of partisan online news and phenomena such as echo chambers and filter bubbles remain concerns (Cinelli et al., 2021; Bakshy et al., 2015). To address these issues, researchers have investigated the political biases embedded in online platforms, including search engines like Google (Robertson et al., 2018) and social media platforms like Facebook (Bakshy et al., 2015), Twitter (Chen et al., 2021), and YouTube (Hosseinmardi et al., 2021). Other relevant research has focused on how users interact with different information sources and their consumption patterns (Garimella et al., 2021; Robertson et al., 2023; González-Bailón et al., 2023). Such analyses generally involve assessing the political leanings of numerous domains, but such datasets have been rare and often lack comprehensive coverage (see discussion in the Technical Validation section).
In this study, we employ DomainDemo-univariate and Eq. (3) to create data-driven metrics that assess the audience partisanship of domains by examining the party affiliation of the users who share these domains. It is important to note that our audience-based metrics do not evaluate the content characteristics of these domains. However, previous research indicates that audience characteristics are closely associated with the leanings of these domains (Gentzkow and Shapiro, 2010; Buntain et al., 2023).
We offer two versions of the audience partisanship scores: one based on inferred user partisanship and another based on party registration. Details of these metrics are provided in Table 3. The audience partisanship scores based on party registration cover fewer domains than other metrics due to the large number of missing values in the party registration column.
Other metrics
We also release three additional audience-based metrics: age deviation, race deviation, and gender leaning. These metrics can be used to understand the sharing patterns conditioned on these demographic variables. Details of the metrics are provided in Table 3.
Examples
Figure 1: Distributions of the derived metrics for all domains, along with detailed information for three domains cnn.com, news9.com, and wickedlocal.com. The left column presents the joint distributions of our derived metrics and the unique number of users who share each domain across the entire dataset. The color coding represents the number of domains within each grid cell. The symbols indicate the locations of the three domains. The three columns on the right provide detailed distributions of users in various demographic dimensions for the three domains, respectively. Sub-figures (b-d) highlight the discrepancies between the observed user distribution and the baseline distribution across U.S. states for the three domains. The color coding indicates the $F_{\delta,s}\log_{2}(F_{\delta,s}/F_{s})$ value in each state. The bar plots display both the baseline distribution and the distribution of users in each demographic category for the domain of interest. The baseline distribution represents the patterns observed across all domains in the dataset.
To aid readers in comprehending the derived metrics, we present the distributions for all domains in the dataset and provide detailed information for three example domains in Figure 1. cnn.com, a national news outlet, has a user base closely aligned with the baseline. Consequently, its localness ($\mathcal{L}_{\delta}=0.013$), race deviation ($\mathcal{L}_{\delta}=0.002$), age deviation ($\mathcal{L}_{\delta}=0.049$), and gender leaning ($\mathcal{P}_{\delta}=-0.033$) scores are near zero. cnn.com is shared more often by Democratic users and less often by Republican users than the baseline, resulting in an audience partisanship score of $-0.132$.
The second example, news9.com, is a local news outlet in Oklahoma City, Oklahoma. It is shared by fewer users than cnn.com and has a localness score of 2.072, indicating a localized audience. Figure 1(c) shows that news9.com is over-represented in Oklahoma, confirming its local nature. Additionally, news9.com is shared more often by Republican users and less often by Democratic users compared to the baseline, leading to an audience partisanship score of 0.297. Its user base has race ($\mathcal{L}_{\delta}=0.051$) and gender ($\mathcal{P}_{\delta}=0.026$) profiles similar to the baseline but is shared more often by older users ($\mathcal{L}_{\delta}=0.086$).
The third example is wickedlocal.com, a local news source in Boston, Massachusetts. Figure 1(d) indicates that it is over-represented in Massachusetts, consistent with its localness score of 2.221. Unlike news9.com, wickedlocal.com is shared more often by Democratic users ($\mathcal{P}_{\delta}=-0.387$) and even more often by older users ($\mathcal{L}_{\delta}=0.246$). Otherwise, the user base of wickedlocal.com has a similar profile in terms of race ($\mathcal{L}_{\delta}=0.071$) and gender ($\mathcal{P}_{\delta}=0.060$) to that of news9.com.
Due to space constraints, we can only provide detailed demographic distributions for these three example domains. We have released an interactive app to allow readers to explore the patterns of other domains in our dataset at domaindemoexplorer.streamlit.app.
Technical Validation
In this section, we demonstrate the robustness of our dataset by first validating the demographic and partisanship measures from the voter records. We then compare the localness and audience partisanship scores of news domains against existing classifications.
Agreement with other data sources
In previous studies, our voter file measure of inferred partisanship was validated at the county level, and the race/ethnicity measure was compared against another voter file-based inference, yielding positive results (Shugars et al., 2021). Another investigation found that the demographic composition of our panel closely resembled representative survey samples of U.S. registered voters on Twitter (Hughes et al., 2021). At the individual level, the voter file measures were validated for 182 panel members also present in the survey samples, showing good matching for gender and reasonable matching for partisanship, but more significant mismatches for race/ethnicity minorities (Hughes et al., 2021). Overall, these findings demonstrate the general validity of the Twitter panel measures, and the domain-sharing statistics from our dataset can be considered generally representative of the sharing patterns of U.S. registered voters on Twitter.
However, the panel matching approach used in this work is not without risks. One source of error arises from the biases inherent in voter files. For instance, there are documented race/ethnicity disparities in who is listed in compiled voter records such as the TargetSmart data we use (Jackman and Spahn, 2021). Additionally, our race/ethnicity measure is inferred based on voter records for most states, implying a potential risk for biases in the categorization of individuals into race/ethnicity groups. Another potential source of error is the linkage process to Twitter users, where individuals with rare names are more likely to be included in the Twitter panel. In terms of representativity, the Twitter panel overrepresents white users and notably underrepresents Hispanics and Asians compared to the survey samples, and it also includes more females (Hughes et al., 2021). Therefore, our DomainDemo-univariate and DomainDemo-multivariate datasets will provide lower percentages of shares by racial minorities and men than a dataset fully representative of U.S. registered voters on Twitter would.
Domain localness metric
Table 4: Summary of existing classifications of local and national news outlets. From left to right, we report the dataset, the number of local domains, the number of national domains, the total number of domains, and the description of the dataset.
Dataset Local National All Description
Cronin et al. (2023) 1,390 276 1,666 Cronin et al. (2023) present a manually coded list of news outlets categorized as local, national, or international. We exclude international news sources from our study.
Fischer et al. (2020) 6,455 511 6,966 Fischer et al. (2020) offer a classification of news domains into categories such as local, regional, national, international, and technical. Their labels are derived from multiple sources, including Yin (2018) and manual annotation. For our analysis, we retain only the local and national domains. The authors do not provide specific criteria to distinguish regional from local categories and often combine them in their analysis. Therefore, we treat all regional domains as local in our study.
Horne et al. (2022) 284 0 284 Horne et al. (2022) compile a list of local news domains sourced from 50states.com/news.
Yin (2018) 6,493 0 6,493 Yin (2018) compile a list of local news outlets. The list is assembled from multiple sources, such as usnpl.com and stationindex.com.
abyz 10,124 290 10,414 abyznewslinks.com provides a list of newspapers and news media in the U.S., classifying them as either local or national. We obtained the data from its website.
In Figure 1, we present the discrepancies between the observed user distribution and the baseline distribution, i.e., $F_{\delta,s}\log_{2}(F_{\delta,s}/F_{s})$, across different states for three domains. The results indicate that our localness metric can effectively capture and quantify the audience patterns of these domains. To systematically validate our localness score, we focus on news media and compile five existing classifications of local and national news outlets. Table 4 provides a summary of the statistics and information for these datasets. While these lists primarily classify news outlets based on coverage and production perspectives, our approach emphasizes the audience perspective.
We merge all these datasets into a single dataset called meta-ln, which contains 12,905 unique domains. Domains are labeled as local or national when there is a consensus among the original sources. Only 40 domains (0.31%) have inconsistent classifications across different datasets. These inconsistencies mainly arise from the varying definitions adopted by different authors for some borderline outlets. For instance, abc7.com is labeled as national by Cronin et al. (2023) but as local by other datasets. We exclude these domains from our analysis and only keep the 4,853 news domains that are present in our dataset for further comparison.
We utilize the Area Under the Receiver Operating Characteristic Curve (AUC) score to assess the alignment between $\mathcal{L}_{\delta}$ and the existing labels. The AUC score essentially measures the probability that our metric assigns higher $\mathcal{L}_{\delta}$ values to local domains compared to national domains (as identified by meta-ln). An AUC score of 0.5 indicates random classification, while a score of 1.0 signifies perfect ranking by $\mathcal{L}_{\delta}$. In our analysis, $\mathcal{L}_{\delta}$ achieves an AUC score of 0.983, indicating minimal discrepancies between meta-ln and $\mathcal{L}_{\delta}$. Although our metric captures different signals than meta-ln, the high agreement level validates the accuracy of our localness metric and the robustness of DomainDemo.
In addition to meta-ln, we compare our localness metric with that of Le Quéré et al. (2022). They also adopt a data-driven approach to quantify the localness of news domains from the audience perspective. Specifically, they quantify the “population reach” of news domains by measuring the distance between the locations of the outlets and the users following the outlets on Twitter while accounting for the population density. Since the population reach metric is continuous, we directly calculate its correlation with $\mathcal{L}_{\delta}$. The intersection between the list shared by Le Quéré et al. (2022) and ours has 1,342 domains and yields a Spearman correlation coefficient of 0.441 ($p<0.001$), suggesting a moderate agreement between the two metrics.
Although our focus here is on news outlets, our dataset encompasses a much wider variety of websites. For instance, our dataset includes information on news domains lacking clear editorial guidelines, such as misinformation websites and “pink slime” websites (Moore et al., 2023). It also covers non-news websites, including government, organizational, entertainment, and e-commerce sites.
Although $\mathcal{L}_{\delta}$ as a continuous value can capture nuanced differences between domains, dichotomizing the value can be beneficial in certain contexts. For news domains, we can use meta-ln to establish a reasonable threshold. By adjusting the threshold value for $\mathcal{L}_{\delta}$, we can compute the corresponding F1 score, which quantifies the agreement between $\mathcal{L}_{\delta}$ and the labels in meta-ln, and identify the optimal threshold that minimizes false positives and false negatives. Our calculations indicate that a threshold of 0.243 yields the highest F1 score of 0.978. When dealing with domains outside of meta-ln, researchers can first annotate a set of domains as local or national using meta-ln, and then use these labels to determine the optimal threshold of $\mathcal{L}_{\delta}$ for their specific study.
Domain audience partisanship metric
Table 5: Summary of the existing domain political leaning scores and their correlations with our audience partisanship scores. From left to right, we report the reference dataset, description of the dataset, number of overlapping domains, and Spearman correlation coefficients with our audience partisanship scores based on inferred partisanship (party) and party registration (party reg) information. All the correlation coefficients are statistically significant at the 0.001 level.
Dataset Description N Party Party reg
Bakshy et al. (2015) Audience-based scores crafted from Facebook data. 398 0.940 0.929
Eady et al. (2020) Media ideology scores based on Twitter data. The authors jointly estimate the idealogy of politicians, users and news sources through the news sharing behaviors on Twitter. 179 0.929 0.929
Buntain et al. (2023) Audience-based scores derived from the Facebook URL dataset. 2,480 0.916 0.898
MBFC (mediabiasfactcheck.com) MBFC (Media Bias Fact Check) provide rater-based political leaning categories for various news domains. We map their labels “far-left,” “left,” “center-left,” “center,” “center-right,” “right,” and “far-right” to numerical values $-1$, $-0.66$, $-0.33$, 00, $0.33$, $0.66$, $1$ for analysis. 2,986 0.765 0.774
Allsides (allsides.com/media-bias/media-bias-ratings) Allsides produces domain bias scores based on their own algorithm. We map their labels “left,” “left-lean,” “center,” “right-lean,” and “right” to numerical values $-1$, $-0.5$, 00, $0.5$, $1$ for analysis. 189 0.736 0.743
Allsides community Similar to the Allsides algorithmic scores, but based on crowdsourced ratings. 189 0.613 0.611
Mturk (Robertson et al., 2018) Crowdsourced ratings from Mturk. 358 0.486 0.491
In a manner similar to the localness metric, we validate the audience partisanship metric by comparing it with the existing classification of domain political leaning, as detailed in Table 5. The table presents the number of domains common to both the reference dataset and our dataset, along with the Spearman correlation coefficients, all of which are positive and statistically significant at the 0.001 level. Our ratings show a high correlation with those of Bakshy et al. (2015), Eady et al. (2020), and Buntain et al. (2023), which are all audience-based scores derived from social media data. The correlation between our metric and other existing political leaning scores that focus on the sources themselves, such as Allsides and MBFC scores, is lower, suggesting that our metric captures different signals. Nonetheless, these findings demonstrate that our metric effectively captures the audience partisanship of various domains.
As detailed in Table 3, we also offer a version of the audience partisanship metric derived from party registration information. This metric shows a strong correlation with the audience partisanship scores based on inferred partisanship, exhibiting a Spearman correlation coefficient of 0.917 ($p<0.001$) across 129,041 overlapping domains. In Table 5, we further compare this metric with existing political leaning scores. As expected, it produces similar results to the audience partisanship scores based on inferred partisanship.
It is important to recognize that obtaining political leaning ratings for extensive sets of domains is challenging. Most existing datasets listed in Table 5 cover only a few hundred domains, with a couple of them covering over 2,000 domains, primarily focusing on news outlets. In contrast, our dataset includes scores for over 129,000 domains, encompassing a diverse array of websites beyond news sources.
By definition, our audience partisanship metric only encodes relative differences, meaning that the zero point does not necessarily indicate politically neutral. Users seeking a binary classification could consider identifying the least biased domains in their contexts and use them to calibrate our metric (Chen et al., 2021).
Data and code availability
Given the sensitive nature of the information about Twitter users in our datasets, we have implemented layered access controls. Since DomainDemo-multivariate can potentially reveal the identities of Twitter users in the dataset when combined with other datasets, it will be hosted within a secure platform. To access this dataset, researchers must complete an application process and sign a data use agreement that prohibits the identification of individual Twitter users. The highly aggregated nature of DomainDemo-univariate reduces the risk of re-identification. Thus, after completing an application process, researchers will be able to download these data files. The derived metrics of the domains, such as localness and audience partisanship scores, are made publicly available.
The details on how to access DomainDemo-multivariate and DomainDemo-univariate will be released soon. At this time, we have released the derived metrics of the domains, together with the code to compute them at github.com/LazerLab/DomainDemo. We also provide an app that allows exploring the derived metrics interactively at domaindemoexplorer.streamlit.app.
Acknowledgements
The authors thank Lisa Friedland for assisting with the data collection and Ronald Robertson for providing the validation data.